OpenStudy (anonymous):

Solve: u^2 + 2u - 3 = 0

OpenStudy (anonymous):

u^2 + 2u - 3 = 0 you can factor this (u+3)(u-1)=0 (u+3)=0 and (u-1)=0 u=-3 and u=1